when you know the heighth of a rectangle is 12 and the width is 12 how do you get the diagnal measerment?
Answer by Nate(3500)

You make a rectangle. Lable the sides 12 and 12 as the height is 12 as well as the width. Draw a diagonal line.
Hmmmm .... does this look familiar? It's a right triangle.
c^2 or the diagonal = a^2 + b^2 or the sides
c^2 = 12^2 + 12^2
c^2 = 288
c = sqrt(288) = sqrt(2*144) = 12sqrt(2)